OpenStudy (anonymous):

simplify: ((x^6y^3)^-1/3)/(x^4y^2)^-1/2

OpenStudy (anonymous):

\[\left( x ^{6}y ^{3} \right)^{-1/3}\div \left( x ^{4}y ^{2} \right)^{-1/2}\]

OpenStudy (kira_yamato):

I've posted all my solutions....

jhonyy9 (jhonyy9):

\[(1/x ^{2y})/(1/x ^{2}y)=(1/x ^{2}y) * x ^{2}y=1\]